What will be the volume of a 3.0M solution using 1.0g of NaCl?
1
Expert's answer
2020-05-10T09:19:43-0400

The molarity of the solution "c" (3.0 M, or mol/L) is defined as:

"c = \frac{n}{V}" ,

where "n" is the number of the moles of the solute (NaCl) and "V" is the volume of the solution. Therefore, the volume of the solution can be expressed from this formulae as:

"V = \frac{n}{c}" .

Let's calculate the number of the moles in 1.0 g of NaCl. The number of the moles can be calculated from the ratio of the mass of NaCl ("m=" 1.0 g) and the molar mass of NaCl ("M =" 58.44 g/mol).

"n = m\/M = 1.0\/58.44 = 0.017" mol

Finally, the volume of the solution:

"V = \frac{0.017}{3.0} = 0.0057" L, or 5.7 mL.

Answer: the volume of a 3.0M solution using 1.0g of NaCl will be 5.7 mL.